OpenStudy (macaylam):

Which system of equations can be graphed to find the solution(s) to 4x^2 = x^2 – 7?

OpenStudy (danjs):

Solutions to 4x^2 = x^2 - 7 Treat each side as its own function, f(x)=4x^2 and g(x)=x^2-7 When those are equal if possible f(x)=g(x) , then there is an intersection point between the two graphs